OpenStudy (anonymous):

Simplify and determine the domain. 1 2 --- - ---- x+6 x-3

OpenStudy (anonymous):

Please Help!

OpenStudy (anonymous):

i thought i was soppose to find a common denomator

sam (.sam.):

Common denominator \[\frac{1}{(x+6)} \times \frac{x-3}{x-3}-\frac{2}{x-3} \times \frac{x+6}{x+6}\]

sam (.sam.):

It is